SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (June 28, 2005)

Attendees: Prafulla Behera, Kris Chaisanguanthum, Yaw Ming Chia, Fang Fang, Selina Li, Brian Lindquist, Dan Walker.

Operations

  • After PEPII recovered from vacuum leak last week, BaBar finally started taking data on Saturday day shift until early Tusday morning when IR6 had a power outage. The DCZ ROM had some trouble due to lost of PEP timing, see the detail under "DCT/DCZ" below.
  • The monthly trigger status report is tomorrow. See the link here.

EMT

  • Dan is modifying the relevant packages for EMT prescaling and testing with createFileTable.
  • Last week, Al suggested to look at timing vs energy to investigate why there were a few entries below the threshold (See some plots of EMT turn-on for muon-pairs from last week). Dan showed these new timing vs energy plots for the E, G, and M.
  • EMT ROM lib was updated to the latest online release last Wednesday when it was installed at IR2. The EMT calibration taken was ok, so were the data.
  • Dan mentioned that the mis-match for the UPC cables between EMC and EMT still happens in EMT calibration as seen in the logfile. This has been happening ever since Novermber 2004 or so. It was believed that this problem only occur in calibration, and doesn't affect data.
  • The .ps file wasn't produced at the end of the EMT calibration, not working probably due to the migration from paw to root in the script.

DCT/DCZ

  • One of the trigger crates in the test stand had overcurrent problem, and it was disassembled to allow Ray to fix the problem. See HN posted for this work. The DCH group will be using the test stand to test the new DCH electronics which needs to have the trigger side working too. They will use the test stand for testing from now on.
  • On Saturday morning, the DCZ crates had to be power cycled.
  • On Tusday morning, the DCZ crates had to be power cycled twice to come back alive. The DCZ ROM didn't see the BLT/ZPD boards after the first power cycle.

L3

  • L3 node trim on Sunday night due to data flow error. This problem actually has been going on since the beginning of the experiment. Currently, there is no other useful info in the core dump about this error. Jim Hamilton will help extract these events to allow studies of this type of errors for node trim. Perhaps the Dataflow group need to look into this too?
  • The L3 node trim due to L3 tracking is still under investigation. See L3 meeting for more detailed discussions.

Data Quality and Monitoring

  • Brian (from Stanford University) is helping us with some of the fast monitoring tasks this summer. He starts working this week.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.